karlieeeeeee:

Is (1, 6) a solution to this system of equations? 4x + 2y = 16 8x + y = 14

Convert:

rtgijorojtgjiojoiyjiohjioiyj did you not see what I said-

Convert:

Try this one by yourself because I helped you with the other ones and this is the same kind of question- :)

karlieeeeeee:

: (

Convert:

C'mon what did I do for your last question? If you need to look back here they are: https://questioncove.com/study#/updates/603e5816b7c024e5dbafb00 https://questioncove.com/study#/updates/603e572d19b8e28c57436700

CripQUEZZ:

Jus use the same method I showed you last time once u get it you’ll start to understand it

CripQUEZZ:

What there saying is that X = 1 And Y = 6 Now plug those in to the equation like 4(1) + 2(6) = 16 So do the multiplication (4 x 1) + (2 x 6) = 16 Then once u do that do the same to the second equation and youll get your answer

Convert:

Yeah, like they said you take the 1st point (x) aka 1 and the second point (y) aka 6 and then plug them into the x and y of the equation, 4(1) + 2(6) = 16 and if that's true and it does equal 16, then it would be a solution. Do you get what me and @CripQUEZZ did?

CripQUEZZ:

It’s pretty easy once you get the hang of it once you start doing more of these you’ll get in the rhythm of it

CripQUEZZ:

Jus always know for these equations you take the two numbers the first one is gonna be the x and the second number will be y and then you jus plug it in to your answer by multiplying them and you’ll get the answer and you can do the same for the second equation

karlieeeeeee:

Oh okay thanks bae

CripQUEZZ:

Np